A plane wave has a equation Y=25sin(120t-40x),Find the
a- wavelength,
b- wave velocity,
c -frequency of the wave,
d - period of the wave,
e -amplitude of the wave .
(where y and x are in meters,t is in seconds )
1
Expert's answer
2020-07-27T09:35:51-0400

"\xi=A\sin(\omega t-kx+\phi_0)"


"y=25\sin(120 t-40x)"


We have


a) "k=2\pi/\lambda\to \lambda=2\pi/k=2\cdot3.14/40=0.157(m)"


b) "k=\omega/v\to v=\omega/k=120/40=3(m/s)" (phase velocity)


c) "\omega=2\pi\nu\to\nu=\omega/(2\pi)=120/(2\cdot3.14)=19.1(Hz)"


d) "T=1/\nu=1/19.1=0.052(s)"


e) "A=25(m)"